OpenStudy (anonymous):

Lyle has a ten-cup box of cereal. If one serving is 3/4 cup, how many servings are in the box?

OpenStudy (anonymous):

10/(3/4) = 40/3 =~ 13.3333 so a whole number answer would be 13 full servings

OpenStudy (anonymous):

Thank U